OpenStudy (anonymous):

If f(x) = log2 (x + 4), what is f^(−1)(3)?

OpenStudy (isaiah.feynman):

Find the inverse first, then substitute x for 3.

OpenStudy (anonymous):

inverse of what?

OpenStudy (isaiah.feynman):

Inverse of f(x), it says so in the question.

OpenStudy (anonymous):

so what would be the inverse of x here

OpenStudy (isaiah.feynman):

Not the inverse of x, the inverse of the function!

OpenStudy (isaiah.feynman):

To find the inverse of a function, use the following procedure! change f(x) to y switch the positions of y and x solve for y the new function is the inverse.

OpenStudy (anonymous):

where is there a y...

OpenStudy (agent0smith):

There isn't a y. He said change f(x) to y. Like this f(x) = log2 (x + 4), y=log2 (x+4)